What is the sum of the finite arithmetic series?
(-5)+0+5+10+...+65
a. 900
b. 455
c. 450
d. 445

Answer by richwmiller(17219)

65 = -5 + (n - 1)5
n=15
S = ½(-5 + 65)15=450
also
S = ½(2a + d(n-1))n
S = ½(-10 + 5(14))15
S=1/2(60)15=450
c. 450